# Bayes Adaptive Monte Carlo Tree Search for Offline Model-based Reinforcement Learning

### v1

@@ Appendix E Computation Cost of Sampled EfficientZero on D4RL MuJoCo
Table 6: Training time (in hours) of Sampled EfficientZero for each evaluation task. The mean and standard deviation from repeated experiments are reported.
Appendix F Ablation Study on the Reward Penalty
@@
random HalfCheetah 29.20 $\pm$ 2.00 36.23 $\pm$ 1.04 32.76 $\pm$ 1.16 34.80 $\pm$ 1.39 38.78 $\pm$ 1.65 39.64 $\pm$ 2.86
random Hopper 33.83 $\pm$ 0.10 31.56 $\pm$ 0.12 31.47 $\pm$ 0.03 9.16 $\pm$ 0.16 7.44 $\pm$ 0.14 6.97 $\pm$ 0.07
random Walker2d 21.89 $\pm$ 0.07 21.59 $\pm$ 0.32 21.45 $\pm$ 0.53 17.53 $\pm$ 6.16 21.53 $\pm$ 0.42 21.41 $\pm$ 0.64
medium HalfCheetah 70.47 $\pm$ 3.52 75.84 $\pm$ 3.81 56.54 $\pm$ 5.20 61.64 $\pm$ 4.58 60.84 $\pm$ 2.00 41.49 $\pm$ 2.29
medium Hopper 97.75 $\pm$ 7.09 96.70 $\pm$ 14.0 98.25 $\pm$ 3.42 102.8 $\pm$ 2.29 104.4 $\pm$ 1.88 93.68 $\pm$ 11.4
medium Walker2d 82.24 $\pm$ 1.85 74.73 $\pm$ 3.25 75.41 $\pm$ 4.17 82.61 $\pm$ 0.86 57.01 $\pm$ 7.24 57.97 $\pm$ 15.4
med-replay HalfCheetah 61.16 $\pm$ 1.60 65.45 $\pm$ 0.81 62.50 $\pm$ 0.18 42.10 $\pm$ 2.85 36.65 $\pm$ 2.39 44.03 $\pm$ 7.35
med-replay Hopper 106.3 $\pm$ 0.13 101.8 $\pm$ 3.46 93.91 $\pm$ 4.25 107.9 $\pm$ 0.07 84.11 $\pm$ 2.97 91.81 $\pm$ 11.5
med-replay Walker2d 92.13 $\pm$ 5.13 95.06 $\pm$ 2.11 97.54$\pm$ 1.93 88.61 $\pm$ 5.21 97.33 $\pm$ 3.51 98.19 $\pm$ 1.23
med-expert HalfCheetah 80.53 $\pm$ 6.63 76.16 $\pm$ 10.3 90.52 $\pm$ 4.13 51.76 $\pm$ 5.31 26.60 $\pm$ 1.46 29.88 $\pm$ 2.28
med-expert Hopper 112.2 $\pm$ 0.29 108.3 $\pm$ 0.22 107.8 $\pm$ 0.37 106.8 $\pm$ 6.34 81.76 $\pm$ 6.45 86.79 $\pm$ 18.7
med-expert Walker2d 107.7 $\pm$ 0.82 110.0 $\pm$ 1.74 84.71 $\pm$ 0.87 110.8 $\pm$ 1.72 110.2 $\pm$ 0.91 53.35 $\pm$ 38.0
@@
Table 7: Comparison of the proposed algorithms with their corresponding versions without the reward penalty (i.e., $\lambda=0$). The definitions of the values in this table are consistent with those in Table 5.
@@
To demonstrate the necessity of incorporating the reward penalty in offline MBRL, we conduct an ablation study by setting $\lambda$ in Eq. (5) to 0, resulting in ablated versions of our proposed three algorithms. The results are presented in Table 7. First, the average performance of the algorithms with the reward penalty is consistently better, demonstrating the importance of using reward penalties in offline MBRL to prevent the overexploitation of the learned world models (which can be inaccurate). Second, the supervised-learning-based algorithm (i.e., BA-MCTS-SL ($\lambda=0$)) is less affected by the absence of the reward penalty, compared to the policy-gradient-based methods. Notably, BA-MCTS-SL ($\lambda=0$) and BA-MCTS-SL achieve comparable performance in the Hopper and Walker2d tasks. This shows an additional advantage of BA-MCTS-SL – its reduced sensitivity to model inaccuracies. Lastly, there are instances where superior performance is achieved with $\lambda$ set to 0. For example, BA-MCTS-SL ($\lambda=0$) performs better than BA-MCTS-SL in 5 out of 12 tasks. This suggests that the performance of our algorithms in Tables 1 and 5 could be further improved by adjusting hyperparameters such as $\lambda$ [^12] As mentioned in Section 5, we retain most of the hyperparameter settings from Optimized (Lu et al. (2022)) to ensure that the performance improvements are attributed to our algorithm design..
Appendix G Details of the Tokamak Control Tasks
STATE SPACE
ACTION SPACE
Targets Current Target, Density Target
Table 8: The state and action spaces of the tokamak control tasks.
Nuclear fusion is a promising energy source to meet the world’s growing demand. It involves fusing the nuclei of two light atoms, such as hydrogen, to form a heavier nucleus, typically helium, releasing energy in the process. The primary challenge of fusion is confining a plasma, i.e., an ionized gas of hydrogen isotopes, while heating it and increasing its pressure to initiate and sustain fusion reactions. The tokamak is one of the most promising confinement devices. It uses magnetic fields acting on hydrogen atoms that have been ionized (given a charge) so that the magnetic fields can exert a force on the moving particles (Pironti & Walker (2005)).
Char et al. (2024) trained a deep recurrent network as a dynamics model for the DIII-D tokamak, a device located in San Diego, California, and operated by General Atomics, using a large dataset of operational data from that device. A typical shot (i.e., episode) on DIII-D lasts around 6-8 seconds, consisting of a one-second ramp-up phase, a multi-second flat-top phase, and a one-second ramp-down phase. The DIII-D also features several real-time and post-shot diagnostics that measure the magnetic equilibrium and plasma parameters with high temporal resolution. The authors demonstrate that the learned model predicts these measurements for entire shots with remarkable accuracy. Thus, we use this model as a “ground truth" simulator for tokamak control tasks. Specifically, we generate a dataset of 725270 transitions for offline RL and evaluate the learned policy using this data-driven simulator.
The state and action spaces for the tokamak control tasks are outlined in Table 8. For detailed physical explanations of their components, please refer to (Abbate et al. (2021); Char et al. (2023); Ariola et al. (2008)). The state space consists of five scalar values and six profiles which are discretized measurements of physical quantities along the minor radius of the toroid. After applying principal component analysis (Maćkiewicz & Ratajczak (1993)), the pressure profile is reduced to two dimensions, while the other profiles are reduced to four dimensions each. In total, the state space comprises 27 dimensions. The action space includes direct control actuators for neutral beam power, torque, gas, ECH power, current, and magnetic field, as well as target values for plasma density and plasma shape, which are managed through a lower-level control module. Altogether, the action space consists of 14 dimensions. While for certain tasks, it is possible to prune the state and action spaces to reduce the learning complexity, we have chosen not to apply any domain-specific knowledge in these evaluations for general RL algorithms. We reserve the domain-specific applications of our algorithms, which would require more domain knowledge and engineering efforts, as an important future work.
We select a reference shot from DIII-D, which spans 251 time steps, and use its trajectories of Ion Rotation, Electron Temperature, and $\beta_{n}$ as targets for three tracking tasks. Specifically, $\beta_{n}$ is the normalized ratio between plasma pressure and magnetic pressure, a key quantity serving as a rough economic indicator of efficiency. Since the tracking targets vary over time, we include the time step as part of the policy input. The reward function for each task is defined as the negative squared tracking error of the corresponding component (i.e., temperature, rotation, or $\beta_{n}$) at each time step, and the reward is normalized by the episode horizon (i.e., 251 time steps). Notably, for policy learning, the reward function is provided rather than learned from the offline dataset as in D4RL tasks; and the dataset does not include the reference shot or any nearby, similar shots.

### v4

@@ F.4 Comparison with MuZero and its Variants
Table 10: Training time (in hours) of our algorithms and Sampled EfficientZero for each evaluation task. Results are presented as the mean and standard deviation from three repeated experiments.
The performance gap may be attributed to the challenges of applying MuZero’s design to the offline setting. World model learning is particularly difficult when the state-action space is vast but training data is limited. Sampled EfficientZero integrates model learning and policy training into a single stage, which can significantly increase the learning difficulty compared to our decoupled approach.
F.5 Ablation Study on the Enhancement from Belief Adaptation and MCTS-based Planning
Data Type Environment Optimized BA-MBRL BA-MCTS BA-MCTS-SL
random HalfCheetah 31.7 32.76 (1.16) 36.23 (1.04) 29.20 (2.00)
random Hopper 12.1 31.47 (0.03) 31.56 (0.12) 33.83 (0.10)
random Walker2d 21.7 21.45 (0.53) 21.59 (0.32) 21.89 (0.07)
medium HalfCheetah 45.7 56.54 (5.20) 75.84 (3.81) 70.47 (3.52)
medium Hopper 69.3 98.25 (3.42) 96.70 (14.0) 97.75 (7.09)
medium Walker2d 79.7 75.41 (4.17) 74.73 (3.25) 82.84 (1.85)
med-replay HalfCheetah 58.0 62.50 (0.18) 65.45 (0.81) 61.16 (1.60)
med-replay Hopper 90.8 93.91 (4.25) 101.8 (3.46) 106.3 (0.13)
med-replay Walker2d 65.8 97.54 (1.93) 95.06 (2.11) 92.13 (5.13)
med-expert HalfCheetah 104.2 90.52 (4.13) 76.16 (10.3) 80.53 (6.63)
med-expert Hopper 105.8 107.8 (0.37) 108.3 (0.22) 112.2 (0.29)
med-expert Walker2d 97.1 84.71 (0.87) 110.0 (1.74) 107.7 (0.82)
Average Score 65.2 71.06 74.45 74.62
Table 11: Performance comparison among our algorithm: BA-MCTS, its ablated versions: BA-MBRL and BA-MCTS-SL, as well as the baseline: Optimized (Lu et al., 2022). To ensure a principled ablation study, we reimplement BA-MCTS and its ablated variants using the codebase of Optimized, which systematically investigates design choices in offline MBRL. We make only minimal modifications to the code and hyperparameter settings. This approach ensures that any observed performance improvements are attributable to algorithmic design, rather than implementation differences or tuning tricks.
For a principled ablation study, we reimplement BA-MCTS and its ablated variants based on the codebase of Optimized (Lu et al., 2022), which thoroughly explores design choices in offline MBRL, making minimal changes to the code and hyperparameter settings (see Appendix D). The comparison results are presented in Table 11555The performance of BA-MCTS decreases, relative to the results presented in Table 2, for two main reasons: (1) The original implementation in Table 2 is based on a different codebase than the one used in Optimized; and (2) for the ablation results shown in Table 11, we adopt the reward penalty design from Optimized to enable exact ablation comparisons, rather than using the one described in Eq. (5)., and the training curves of BA-MCTS and its ablated variants are shown in Figure 3.
Firstly, by comparing the following two methods, any observed performance difference can be directly attributed to the benefits of belief adaptation in a BAMDP formulation.
  • Optimized: A strong, non-adaptive offline model-based RL baseline.
  • BA-MBRL: Our ablation variant, which is a minimal modification of Optimized, differing only by the integration of our BAMDP framework that dynamically adapts the belief over the model ensemble via Bayesian updates (Eq. (4)).
As demonstrated, BA-MBRL outperforms the Optimized baseline on 8 out of 12 tasks and achieves a higher average score, providing direct evidence that Bayesian belief adaptation leads to improved final policies.
Secondly, by comparing the following two methods, any observed performance difference can be directly attributed to the benefits of incorporating MCTS-based planning:
  • BA-MBRL: Our ablation variant that incorporates the BAMDP belief adaptation framework but does not use MCTS. Its policy improvement relies solely on Soft Actor Critic.
  • BA-MCTS: Our final algorithm, which builds upon BA-MBRL by integrating the Continuous BAMCP deep search planner.
This substantial gain highlights the effectiveness of using Continuous BAMCP as a powerful policy improvement operator. While BA-MBRL relies on the implicit policy improvement provided by the temporal difference learning process, BA-MCTS leverages deep lookahead search to explicitly find a stronger, non-parametric policy ($\pi_{\text{ret}}$ returned by Algorithm 1) at each decision point. The actor-critic learner is then trained on the high-quality trajectories generated from this improved policy. This “RL+Search” paradigm, where the learner distills knowledge from a stronger planner, is the primary reason for the observed performance leap. This result strongly supports our claim that the deep search component is crucial for achieving state-of-the-art performance in the offline model-based setting.
(a) hc-med-expert
(b) hc-med-replay
(c) hc-medium
(d) hc-random
(e) hp-med-expert
(f) hp-med-replay
(g) hp-medium
(h) hp-random
(i) wk-med-expert
(j) wk-med-replay
(k) wk-medium
(l) wk-random
Figure 3: Performance of our proposed algorithms on D4RL MuJoCo tasks in the ablation study. The results for HalfCheetah, Hopper, and Walker2d are presented in the three rows, respectively. Each subfigure depicts the change in the undiscounted episodic return as a function of training epochs. Experiments are repeated three times with different random seeds, with the solid line representing the mean and the shaded area indicating the 95% confidence interval. For reference, the expert-level episodic returns for HalfCheetah, Hopper, and Walker2d are 12135, 3234.3, and 4592.3, respectively. Note that the training epochs for each algorithm, as listed in Table 5, have been linearly scaled to 800 for better visualization.
F.6 Comparison of Policy Distillation Mechanisms
This section addresses the research question: How can the search outcomes (i.e., $\pi_{\text{ret}}$) be effectively used for policy improvement? We investigate this by comparing two different mechanisms for distilling the knowledge from the MCTS planner back into the parameterized policy network $\pi$.
  • BA-MCTS: Our primary algorithm, which uses a policy gradient method (i.e., SAC) to update the policy network using trajectories generated by the planner.
  • BA-MCTS-SL: An alternative variant where the policy network is updated via supervised learning (SL) to directly mimic the planner’s output distribution $\pi_{\text{ret}}$ by minimizing a cross-entropy loss.
Performance Comparison: As shown in Table 11, BA-MCTS-SL performs similarly to BA-MCTS in terms of the final average score, validating the effectiveness of both policy update mechanisms. This indicates that both policy gradient and supervised learning are viable methods for distilling the results of the deep search. However, we note that BA-MCTS-SL can be less stable on certain complex tasks; for example, it struggles on the Walker2d environments, where a warm-up training phase (using BA-MBRL) is required to establish a good initial policy for the supervised learning to be effective.
Training Stability and Robustness: On the other hand, a key advantage of the SL-based policy update is its training stability. As evident in the training plots (Figure 3), BA-MCTS-SL often exhibits much smoother learning curves compared to the other two algorithms, indicating greater robustness in model selection for deployment. Furthermore, as shown in our ablation study on the reward penalty (Appendix F.7), we found that the SL-based policy update is less sensitive to the degree of pessimism.
F.7 Ablation Study on the Reward Penalty
@@
random HalfCheetah 29.20 (2.00) 36.23 (1.04) 32.76 (1.16) 34.80 (1.39) 38.78 (1.65) 39.64 (2.86)
random Hopper 33.83 (0.10) 31.56 (0.12) 31.47 (0.03) 9.16 (0.16) 7.44 (0.14) 6.97 (0.07)
random Walker2d 21.89 (0.07) 21.59 (0.32) 21.45 (0.53) 17.53 (6.16) 21.53 (0.42) 21.41 (0.64)
medium HalfCheetah 70.47 (3.52) 75.84 (3.81) 56.54 (5.20) 61.64 (4.58) 60.84 (2.00) 41.49 (2.29)
medium Hopper 97.75 (7.09) 96.70 (14.0) 98.25 (3.42) 102.8 (2.29) 104.4 (1.88) 93.68 (11.4)
medium Walker2d 82.24 (1.85) 74.73 (3.25) 75.41 (4.17) 82.61 (0.86) 57.01 (7.24) 57.97 (15.4)
med-replay HalfCheetah 61.16 (1.60) 65.45 (0.81) 62.50 (0.18) 42.10 (2.85) 36.65 (2.39) 44.03 (7.35)
med-replay Hopper 106.3 (0.13) 101.8 (3.46) 93.91 (4.25) 107.9 (0.07) 84.11 (2.97) 91.81 (11.5)
med-replay Walker2d 92.13 (5.13) 95.06 (2.11) 97.54 (1.93) 88.61 (5.21) 97.33 (3.51) 98.19 (1.23)
med-expert HalfCheetah 80.53 (6.63) 76.16 (10.3) 90.52 (4.13) 51.76 (5.31) 26.60 (1.46) 29.88 (2.28)
med-expert Hopper 112.2 (0.29) 108.3 (0.22) 107.8 (0.37) 106.8 (6.34) 81.76 (6.45) 86.79 (18.7)
med-expert Walker2d 107.7 (0.82) 110.0 (1.74) 84.71 (0.87) 110.8 (1.72) 110.2 (0.91) 53.35 (38.0)
@@
Table 12: Comparison of the proposed algorithms with their corresponding versions without the reward penalty (i.e., $\lambda=0$). The definitions of the scores in this table are consistent with those in Table 9.
@@
To demonstrate the necessity of incorporating the reward penalty in offline MBRL, we conduct an ablation study by setting $\lambda$ in Eq. (5) to 0, resulting in ablated versions of our proposed three algorithms. The results are presented in Table 12. First, the average performance of the algorithms with the reward penalty is consistently better, demonstrating the importance of using reward penalties in offline MBRL to prevent the overexploitation of the learned world models (which can be inaccurate). Second, the supervised-learning-based algorithm (i.e., BA-MCTS-SL ($\lambda=0$)) is less affected by the absence of the reward penalty, compared to the policy-gradient-based methods. Notably, BA-MCTS-SL ($\lambda=0$) and BA-MCTS-SL achieve comparable performance in the Hopper and Walker2d tasks. This shows an additional advantage of BA-MCTS-SL – its reduced sensitivity to model inaccuracies. Lastly, there are instances where superior performance is achieved with $\lambda$ set to 0. For example, BA-MCTS-SL ($\lambda=0$) performs better than BA-MCTS-SL in 5 out of 12 tasks. This suggests that the performance of our algorithms in Tables 11 could be further improved by adjusting hyperparameters such as $\lambda$.
Appendix G Statement on the Use of Large Language Models
In the spirit of transparency and in accordance with the ICLR disclosure policy, we detail the use of Large Language Models (LLMs) in the preparation of this manuscript.
LLMs were utilized as an assistive tool in two main capacities: language refinement and technical support during the formalization of proofs. The specific roles were:
  • Language and Readability Enhancement: We employed an LLM to polish the manuscript’s language. This included improving grammar, rephrasing sentences for better clarity, and ensuring a consistent and professional tone throughout the paper.
  • Assistance with Theoretical Proofs: An LLM provided support in the technical writing and verification of our theoretical proofs. This assistance was twofold:
    1. 1.
      Generating source code for complex mathematical equations.
    2. 2.
      Acting as a verification aid by reviewing the logical steps in our derivations to help identify potential inconsistencies or errors.
It is crucial to state that the LLM did not contribute to the original ideation, conceptual development, or the core derivations of the proofs themselves. The intellectual ownership and the scientific reasoning behind all theoretical results remain entirely with the authors. The authors have meticulously reviewed, validated, and take full and final responsibility for all content, including the correctness and rigor of the proofs that were checked with LLM assistance.